OpenStudy (anonymous):

find the largest of 3 consecutive integers such that 3 times the sum of the first and the third integer is equal to 30 more than 4 times the 2nd integer............PLEASE HURRY FIRST RESPONSE GETS AN MEDAL !!!!!

OpenStudy (anonymous):

1st integer = n 2nd = n+1 3rd=n+2 form its equation: 3[(n) + (n+2)] = 30 + 4(n+1) solve now.

OpenStudy (anonymous):

n = 14

OpenStudy (anonymous):

now.. largest integer = n+ 2 = 14 + 2 = 16 ans

OpenStudy (texaschic101):

3(2n + 2) = 30 + 4n + 4 6n + 6 = 34 + 4n 6n - 4n = 34 - 6 2n = 28 n = 14 n + 2 = 14 + 2 = 16 you are correct rukshar...good job explaining :)

OpenStudy (anonymous):

thanks..

OpenStudy (texaschic101):

just telling the truth :)